A group of fun-loving people have decided to play a practical joke on one of their friends, but they can't decide which friend, Alice (A), Betty (B), or Connie (C). Their preferences are:

(BCA)

(ACB)

(CAB)

10

4

5


Who wins the election using the Hare method? Does this violate the Condorcet criterion?

a.

Betty, No

b.

Connie, Yes

c.

Alice, Yes

d.

Connie, No

e.

Betty, Yes
Answer #1

Total votes = 10+4+5=19 and Total seats is 1.

Hare quota is Total votes/ (Total seats+1)+1 = 19/3=6.667

First preferences for each candidate:

A:4, B:10, C:5

Since B is over the Hare quota, B (Betty) is declared the winner

B is also the Condocert winner as it is preferred by 10 people over the remaining 9. So the Condocert criterion is not violated

Correct option is a) Betty, No